AB&M UV Flood Exposure System is up. Please check the log book to see the power intensity.
We received the Flow Switch for the Trion and now it just needs to be installed. We will update you when the equipment is up.

AB&M UV Flood Exposure System is up. Please check the log book to see the power intensity.
We received the Flow Switch for the Trion and now it just needs to be installed. We will update you when the equipment is up.
The lamp has arrived and we will be installing it today.
UP:
Note: These have not been benchmarked since the power outage so your results may vary.
DOWN:
We have relocated the Booties and Hairnet boxes to the locker room next to the shoe cleaner. Please put on booties after using the shoe cleaner in the locker room in order to keep the gowning area clean. As always please keep all personal item (jackets, hats, bags…) is this room as well. For a locker please see Ryan in the INRF.

The lamp replacement for the AB&M UV Flood Exposure System will be here this Thursday.
Jake will not be in today so there will be no user meeting this week.
The INRF & BION cleanrooms will be closing down at 4pm this Friday due to the campus wide power shutdown. We expect the labs to be back up late Monday or worst case Tuesday depending on how smoothly everything powers back up. We apologize for any inconvenience.
Air Handler Unit (AHU) # 7 will be shutdown tomorrow between 7:00AM – 8:00AM. This will affect the INRF cleanroom as the pressure will be negative during this time, so please adjust your workload accordingly.
We apologize for any inconvenience.
The INRF Store now carries Double Sided Polished 4″ wafers. They are priced at $34ea.
Details: 4″ P/B <1-0-0> 1-10 OHM-CM, 375 ± 25 μm, DSP WAFERS
Facilities will be shutting down Air handler #3 for about 30 mins right now. The make up handler should pick up the slack but please keep an eye out for any changes.